As the HIM director of a small community hospital, you are required to prepare a budget for the department and submit to your executive management team. As part of the budget, the evaluation of salaries is part of the analysis. Three employees are eligible for pay raises during the next fiscal year, which is the calendar year for this organization. Your CFO has informed you that your allotted budget should not exceed 6% of the original base salaries for these 3 employees. Complete the table for the following employees and proposed pay raises for each percentage based on the following data. Also, answer the question, do the proposed increases meet/exceed the 6% allowed by your CFO? Please explain. Please show your work in the cell to evaluate the calculations and result.

1. The Coder/Analyst I is getting a 5% raise at the beginning of the calendar/fiscal year.

2. The ROI Specialist/Analyst II is getting a 4% raise at the beginning of the calendar/fiscal year.

3. The Coding Compliance Specialist/Analyst III is getting 2.5% a raise at the beginning of the calendar/fiscal year.

4. Does your proposed increase meet/exceed the 6% allowed by your CFO? Please explain.

Calculations are based on • 40 hour work week (2080 hours in a year). •

Reminder: 52 weeks in a year

Position or Title Current Base Yearly Salary Proposed Raise New Base Yearly Salary
Coder/Analyst I $24,000
ROI Specialist/Analyst II $26,000
Coding Compliance Specialist/Analyst III $37,000
Total $87,000
Answer #1
A B C=A*B D=A+C
Position or Title Current Base Yearly Salary % of raise Proposed Raise New Base Yearly Salary
Coder/Analyst I      24,000.00 5% 1,200.00              25,200.00
ROI Specialist/Analyst II      26,000.00 4% 1,040.00              27,040.00
Coding Compliance Specialist/Analyst III      37,000.00 2.5%       925.00              37,925.00
Total     87,000.00 3,165.00              90,165.00 E
Maximum raise from base salary 6% F
Maximum raise from base salary        5,220.00 G=E*F
Maximum raise from base salary is $ 5,220.
Proposed from base salary is $ 3,165.
So the proposed increases meets the 6% allowable limit as it is less than the maximum raise.